OpenStudy (anonymous):

list the intercepts and test for symmetry y= 3x/x^2+81

OpenStudy (anonymous):

Parenthesis please... You wrote: \[y= \frac{3x}{x^2}+81\] But I think you probably meant: \[y= \frac{3x}{x^2+81}\] This is why parenthesis matter. The two are totally different graphs!

OpenStudy (anonymous):

sorry yes its the second one

OpenStudy (anonymous):

intercepts are when it crosses the x-axis (y=0) and when it crosses the y-axis (x=0). Find what y is when x is zero first here. f(0) = ? (that means plug in 0 for x if that wasn't clear)

OpenStudy (anonymous):

FYI: The correct way to write the second one would have been: y= 3x/(x^2+81) See why? Multiplication before addition. \[f(0) = \frac{3(0)}{(0)^2+81} = \ ?\]

OpenStudy (anonymous):

y = f(x) I'm just making it into a function here so you can see it easier.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im still a little confused

OpenStudy (anonymous):